OpenStudy (anonymous):

Which of the following is the correct point slope equation for the line that passes through the point (-4,-2) and is parallel to the line given by y=5x+44? A. y-2=(5)x-4 B. y+2=5(x+4) C. y+4=5(x+2) D. y-4=5(x-2)

OpenStudy (anonymous):

Point slope equation: y - y1 = m(x - x1) m= the slope (x1,y1) is any point on your line. Do you know what x1= and y1= ? What about the slope?